# Wall crossing and the Fourier–Mukai transform for Grassmann flops

### v1

@@ 2. Geometric setup
Let $E=\oplus_{j=1}^{n}L_{j}$ and $F=\oplus_{i=1}^{n}M_{i}$ be sums of line bundles over a smooth projective variety $B$. Choose $r<n$, and define the vector bundle $V\to B$ to be the total space of
$\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E).$
@@
This vector bundle comes equipped with a (right) action of $G=\operatorname{Aut}(\mathbb{C}^{r})=\operatorname{GL}_{r}$ given fiberwise by
$(X,Y)\cdot g=(g^{-1}X,Yg).$
Let $X_{+}$ and $X_{-}$ denote the GIT quotients $V\sslash_{\pm}G$ associated to the characters $\det^{+1}$ and $\det^{-1}$ respectively. Then
$X_{+}=$ $\operatorname{tot}(S_{+}\otimes p^{*}F^{\vee})\to\operatorname{Gr}(r,E)$
$X_{-}=$ $\operatorname{tot}(S_{-}\otimes p^{*}E)\to\operatorname{Gr}(r,F^{\vee})$
@@
where $p:\operatorname{Gr}(r,E)\to B$ (resp. $p:\operatorname{Gr}(r,F^{\vee})\to B$) denotes the relative Grassmannian of $r$-planes in $E$ (resp. $F^{\vee}$) and $S_{\pm}$ denotes the rank-$r$ tautological bundle on the associated Grassmann bundle. In other words, the fiber of
@@
$f:X_{-}\dashrightarrow X_{+}$
is a relative version of what is known as a Grassmann flop. We construct a resolution of this map in Section 2.3 following [BLVdB16], and compute the associated Fourier–Mukai transform in $K$-theory.
2.1. Cohomology
Let $(\mathbb{C}^{*})^{n}$ act on $F$ by scaling, such that the $i$th factor scales $M_{i}$. Let a different copy of $(\mathbb{C}^{*})^{n}$ act on $E$ similarly. This induces an action of the torus $\mathbb{T}:=(\mathbb{C}^{*})^{2n}$ on both $X_{+}$ and $X_{-}$. Unless otherwise specified, all cohomology groups are $\mathbb{T}$-equivariant. We define $z_{i}=c_{1}^{\mathbb{T}}(M_{i}^{\vee})$ and $x_{i}=c_{1}^{\mathbb{T}}(L_{i}^{\vee})$.
@@ 2.1. Cohomology
(2.1.1) $v\cdot g=vg.$
The $G$-equivariant vector bundle $R\times V\to V$ induces a vector bundle on both $X_{+}$ and $X_{-}$ which, by abuse of notation, we also denote by $R$. On $X_{-}$ it is (the pullback of) the tautological bundle $S_{-}\to\operatorname{Gr}(r,F^{\vee})$ and on $X_{+}$ it is (the pullback of) the dual of the tautological bundle $S_{+}\to\operatorname{Gr}(r,E)$. Denote the equivariant Chern roots of the dual $R^{\vee}$ by $y_{1},\ldots,y_{r}$.
Let
$\pi_{\pm}:X_{\pm}\to B$
denote the projection. To simplify notation we will denote $\pi_{\pm}^{*}(x_{i})$ (resp. $\pi_{\pm}^{*}(z_{j})$) simply by $x_{i}$ (resp. $z_{j}$) when no confusion will result. The equivariant cohomology of $X_{+}$ (resp. $X_{-}$), denoted $H^{*}_{\mathbb{T}}(X_{\pm})$, is generated as an algebra by elements of
$H^{*}(B)[z_{i},x_{i}]_{i=1}^{n}$
and $\sigma_{1},\ldots,\sigma_{r}$, the elementary symmetric polynomials in $y_{1},\ldots,y_{r}$, subject to the relations:
(2.1.2) $I_{+}=\left\{\left[\frac{\prod_{i=1}^{n}(1-x_{i})}{\prod_{j=1}^{r}(1+y_{j})}\right]_{l}=0\ \middle|\ l>n-r\right\}$
in $H^{*}_{\mathbb{T}}(X_{+})$, and
(2.1.3) $I_{-}=\left\{\left[\frac{\prod_{i=1}^{n}(1+z_{i})}{\prod_{j=1}^{r}(1-y_{j})}\right]_{l}=0\ \middle|\ l>n-r\right\}.$
in $H^{*}_{\mathbb{T}}(X_{-})$. In the above formulas, $\left[-\right]_{l}$ denotes the degree $l$ part. We will denote by $R_{\mathbb{T}}$ the ring $H^{*}_{\mathbb{T}}(pt)$, and by $S_{\mathbb{T}}$ its localization with respect to non-zero homogeneous elements. We will also make use of the completion $\hat{S}_{\mathbb{T}}$, defined by
$\hat{S}_{\mathbb{T}}=\{\sum_{d\in\mathbb{Z}}a_{d}|\text{there exists $d_{0}\in\mathbb{Z}$ such that $a_{d}=0$ for $d<d_{0}$}\},$
@@
The connected components of the $\mathbb{T}$-fixed locus of $X_{+}$ (resp. $X_{-}$) are indexed by subsets of $\{1,\ldots,n\}$ of size $r$, corresponding to choosing an $r\times r$ minor of $\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$ (resp. $\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})$). Thus each connected component is isomorphic to $B$. Let $D_{+}$ (resp. $D_{-}$) denote the set indexing these fixed loci. We may represent an element of $D_{\pm}$ by a function $\delta:\{1,\ldots,r\}\to\{1,\ldots,n\}$ such that $\delta(i)<\delta(j)$ for all $i<j$. We will further denote $\delta(i)$ by $\delta_{i}$.
@@
  1. (1)
    For $\delta^{-}\in D_{-}$, let $B_{\delta^{-}}$ denote the associated fixed locus, isomorphic to $B$. Then $R|_{B_{\delta^{-}}}\cong\oplus_{i=1}^{r}M_{\delta^{-}_{i}}^{\vee}$.
  2. (2)
    For $\delta^{+}\in D_{+}$, let $B_{\delta^{+}}$ denote the associated fixed locus, isomorphic to $B$. Then $R|_{B_{\delta^{+}}}\cong\oplus_{j=1}^{r}L_{\delta^{+}_{j}}^{\vee}$.
@@
$\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})\sslash_{\theta_{-}}G.$
There is a tautological map of vector bundles $q^{*}(\oplus_{i=1}^{r}M_{\delta^{-}_{i}})\to R^{\vee}$ over the stack
$[\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})/G,]$
where $q:[\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})/G,]\to B$ is the projection. The $\theta_{-}$-stable locus of $\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})$ is exactly the locus where this map is an isomorphism. ∎
@@
  2. (2)
    In $H^{*}_{\mathbb{T}}(X_{+})$, up to a choice of ordering, the restriction of $y_{i}$ to $B_{\delta^{+}}$ is $-x_{\delta^{+}_{i}}$.
2.2. The associated abelian quotients
Let $T\subset G$ denote the maximal torus in $G$ consisting of diagonal matrices. The characters $\det^{\pm}$ restrict to characters of $T$. We define the abelian quotients $X_{T,\pm}$ associated to $X_{\pm}$ to be the GIT quotients $V\sslash_{\pm}T$. They are fiber products over $B$ of $r$ copies of $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)$ and $r$ copies of $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)$, resp.:
$X_{T,{-}}$ $=\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)\times_{B}\cdots\times_{B}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)$
$X_{T,{+}}$ $=\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)\times_{B}\cdots\times_{B}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right).$
@@ 2.2. The associated abelian quotients
$\theta_{j}((t_{1},\ldots,t_{r}))=\prod_{i=1}^{j}t_{i}\cdot\prod_{i=j+1}^{r}t_{i}^{-1},$
@@
$X_{T,j}=\prod_{i=1}^{j}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)\times_{B}\prod_{i=j+1}^{r}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)$
@@
The $\mathbb{T}$-fixed loci of $X_{T,\pm}$ are indexed by (not necessarily injective) functions $f:\{1,\ldots,r\}\to\{1,\ldots,n\}$. For such an $f$, the associated fixed locus $B_{f}$ is the subspace of $X_{T,+}$ such that, on the $i$th factor of $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)$ in the fiber product, the homogeneous coordinates are all zero except for in the $f(i)$th factor of $\mathbb{P}(E)$. In other words,
$B\cong B_{f}=\mathbb{P}(L_{f(1)})\times_{B}\cdots\times_{B}\mathbb{P}(L_{f(r)})\subset\mathbb{P}(E)\times_{B}\cdots\times_{B}\mathbb{P}(E)\subset X_{T,+}.$
@@
We will make use of the following weak form of the abelian/non-abelian correspondence in cohomology:
@@
$H^{*}_{\mathbb{T}}(X_{\pm})\to(H^{*}_{\mathbb{T}}(V^{s,\pm}(G)/T))^{W}$
is an isomorphism, where $V^{s,\pm}(G)$ denotes the stable locus of $V$ with respect to $\det^{\pm}:G\to\mathbb{C}^{*}$ and $W$ denotes the Weyl group. Furthermore, the pullback map
$H^{*}_{\mathbb{T}}(X_{T,\pm})\to H^{*}_{\mathbb{T}}(V^{s,\pm}(G)/T)$
@@
We note that although not originally stated in this context, the result holds equivariantly, as one can check from the explicit presentation of $H^{*}_{\mathbb{T}}(X_{\pm})$ given in Section 2.1. The proposition allows us to lift cohomology classes in $X_{\pm}$ to $W$-invariant classes in $H^{*}_{\mathbb{T}}(X_{T,\pm})$.
For future use we will use $F_{\pm}$ to denote the set of functions $f:\{1,\ldots,r\}\to\{1,\ldots,n\}$, i.e. $F_{\pm}$ indexes the set of fixed loci of $X_{T,\pm}$. We will use $In_{\pm}$ to denote the set of injective functions $\{1,\ldots,r\}\to\{1,\ldots,n\}$, noting that $In_{\pm}$ indexes the fixed loci of $V^{s,\pm}(G)/T.$
2.3. Fourier–Mukai transform
We can define a resolution of the birational map
$f:X_{-}\dashrightarrow X_{+}$
as follows. Let
$\tilde{V}=\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$
and let $\tilde{G}=GL_{r}\times GL_{r}$. Consider the right action of $\tilde{G}$ on $\tilde{V}$ given by
$(X,U,Y)\cdot(g_{1},g_{2})=(g_{1}^{-1}X,g_{2}^{-1}Ug_{1},Yg_{2}).$
Let $\tilde{\theta}\in\hom(\tilde{G},\mathbb{C}^{*})$ denote the character

### v2

@@ 2 Geometric setup
Let $E=\bigoplus_{j=1}^{n}L_{j}$ and $F=\bigoplus_{i=1}^{n}M_{i}$ be sums of line bundles over a smooth projective variety $B$. Choose $r<n$, and define the vector bundle $V\to B$ to be the total space of
$\mathcal{H}{\rm om}(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E).$
@@
This vector bundle comes equipped with a (right) action of $G=\operatorname{Aut}(\mathbb{C}^{r})={\rm GL}_{r}$ given fiberwise by
$(X,Y)\cdot g=\bigl{(}g^{-1}X,Yg\bigr{)}.$
Let $X_{+}$ and $X_{-}$ denote the GIT quotients $V\sslash_{\pm}G$ associated to the characters $\det^{+1}$ and $\det^{-1}$, respectively. Then
$X_{+}=\operatorname{tot}\bigl{(}S_{+}\otimes p^{*}F^{\vee}\bigr{)}\to\operatorname{Gr}(r,E),\qquad X_{-}=\operatorname{tot}(S_{-}\otimes p^{*}E)\to\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)},$
@@
where $p\colon\operatorname{Gr}(r,E)\to B$ (resp. $p\colon\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}\to B$) denotes the relative Grassmannian of $r$-planes in $E$ (resp. $F^{\vee}$) and $S_{\pm}$ denotes the rank-$r$ tautological bundle on the associated Grassmann bundle. In other words, the fiber of
@@
$f\colon\ X_{-}\dashrightarrow X_{+}$
is a relative version of what is known as a Grassmann flop. We construct a resolution of this map in Section 2.3 following [9], and compute the associated Fourier–Mukai transform in $K$-theory.
2.1 Cohomology
Let $(\mathbb{C}^{*})^{n}$ act on $F$ by scaling, such that the $i$th factor scales $M_{i}$. Let a different copy of $(\mathbb{C}^{*})^{n}$ act on $E$ similarly. This induces an action of the torus $\mathbb{T}:=(\mathbb{C}^{*})^{2n}$ on both $X_{+}$ and $X_{-}$. Unless otherwise specified, all cohomology groups are $\mathbb{T}$-equivariant. We define $z_{i}=c_{1}^{\mathbb{T}}\bigl{(}M_{i}^{\vee}\bigr{)}$ and $x_{i}=c_{1}^{\mathbb{T}}\bigl{(}L_{i}^{\vee}\bigr{)}$.
@@ 2.1 Cohomology
$v\cdot g=vg.$ (2.1)
The $G$-equivariant vector bundle $R\times V\to V$ induces a vector bundle on both $X_{+}$ and $X_{-}$ which, by abuse of notation, we also denote by $R$. On $X_{-}$ it is (the pullback of) the tautological bundle $S_{-}\to\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}$ and on $X_{+}$ it is (the pullback of) the dual of the tautological bundle $S_{+}\to\operatorname{Gr}(r,E)$. Denote the equivariant Chern roots of the dual $R^{\vee}$ by $y_{1},\dots,y_{r}$.
Let $\pi_{\pm}\colon X_{\pm}\to B$ denote the projection. To simplify notation, we will denote $\pi_{\pm}^{*}(x_{i})$ (resp. $\pi_{\pm}^{*}(z_{j})$) simply by $x_{i}$ (resp. $z_{j}$) when no confusion will result. The equivariant cohomology of $X_{+}$ (resp. $X_{-}$), denoted $H^{*}_{\mathbb{T}}(X_{\pm})$, is generated as an algebra by elements of $H^{*}(B)[z_{i},x_{i}]_{i=1}^{n}$ and $\sigma_{1},\dots,\sigma_{r}$, the elementary symmetric polynomials in $y_{1},\dots,y_{r}$, subject to the relations:
$I_{+}=\left\{\left[\frac{\prod_{i=1}^{n}(1-x_{i})}{\prod_{j=1}^{r}(1+y_{j})}\right]_{l}=0\ \middle|\ l>n-r\right\}$
in $H^{*}_{\mathbb{T}}\bigl{(}X_{+}\bigr{)}$, and
$I_{-}=\left\{\left[\frac{\prod_{i=1}^{n}(1+z_{i})}{\prod_{j=1}^{r}(1-y_{j})}\right]_{l}=0\ \middle|\ l>n-r\right\}$
in $H^{*}_{\mathbb{T}}(X_{-})$. (See [20, Example 14.6.6] for an equivalent presentation in the context of Chow rings, the above presentation follows by the same argument.) In the above formulas, $\left[-\right]_{l}$ denotes the degree $l$ part. We will denote by $R_{\mathbb{T}}$ the ring $H^{*}_{\mathbb{T}}(pt)$, and by $S_{\mathbb{T}}$ its localization with respect to non-zero homogeneous elements. We will also make use of the completion $\hat{S}_{\mathbb{T}}$, defined by
$\hat{S}_{\mathbb{T}}=\biggl{\{}\sum_{d\in\mathbb{Z}}a_{d}\,\bigg{|}\,\text{there exists $d_{0}\in\mathbb{Z}$ such that $a_{d}=0$ for $d<d_{0}$}\biggr{\}},$
@@
The connected components of the $\mathbb{T}$-fixed locus of $X_{+}$ (resp. $X_{-}$) are indexed by subsets of $\{1,\dots,n\}$ of size $r$, corresponding to choosing an $r\times r$ minor of $\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$ (resp. $\mathcal{H}{\rm om}(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})$). Thus each connected component is isomorphic to $B$. Let $D_{+}$ (resp. $D_{-}$) denote the set indexing these fixed loci. We may represent an element of $D_{\pm}$ by a function $\delta\colon\{1,\dots,r\}\to\{1,\dots,n\}$ such that $\delta(i)<\delta(j)$ for all $i<j$. We will further denote $\delta(i)$ by $\delta_{i}$.
@@
  1. $1$
    For $\delta^{-}\in D_{-}$, let $B_{\delta^{-}}$ denote the associated fixed locus, isomorphic to $B$. Then
    $R|_{B_{\delta^{-}}}\cong\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}}^{\vee}.$
  2. $2$
    For $\delta^{+}\in D_{+}$, let $B_{\delta^{+}}$ denote the associated fixed locus, isomorphic to $B$. Then
    $R|_{B_{\delta^{+}}}\cong\bigoplus_{j=1}^{r}L_{\delta^{+}_{j}}^{\vee}.$
@@ 2. $2$
$\mathcal{H}{\rm om}\biggl{(}\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r}\biggr{)}\sslash_{\theta_{-}}G.$
There is a tautological map of vector bundles $q^{*}\bigl{(}\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}}\bigr{)}\to R^{\vee}$ over the stack
$\biggl{[}\mathcal{H}{\rm om}\biggl{(}\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r}\biggr{)}/G\biggr{]},$
where $q\colon\bigl{[}\mathcal{H}{\rm om}\bigl{(}\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r}\bigr{)}/G\bigr{]}\to B$ is the projection. The locus where this map is an isomorphism is exactly the $\theta_{-}$-stable locus of $\mathcal{H}{\rm om}\bigl{(}\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r}\bigr{)}$. ∎
@@ 1. $1$
  2. $2$
    In $H^{*}_{\mathbb{T}}\bigl{(}X_{+}\bigr{)}$, up to a choice of ordering, the restriction of $y_{i}$ to $B_{\delta^{+}}$ is $-x_{\delta^{+}_{i}}$.
2.2 The associated abelian quotients
Let $T\subset G$ denote the maximal torus in $G$ consisting of diagonal matrices. The characters $\det^{\pm}$ restrict to characters of $T$. We define the abelian quotients $X_{T,\pm}$ associated to $X_{\pm}$ to be the GIT quotients $V\sslash_{\pm}T$. They are fiber products over $B$ of $r$ copies of $\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\bigr{)}$ and $r$ copies of $\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)}$, respectively:
$X_{T,{-}}=\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\bigr{)}\times_{B}\cdots\times_{B}\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\bigr{)},$
$X_{T,{+}}=\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)}\times_{B}\cdots\times_{B}\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)}.$
@@ 2.2 The associated abelian quotients
$\theta_{j}((t_{1},\dots,t_{r}))=\prod_{i=1}^{j}t_{i}\cdot\prod_{i=j+1}^{r}t_{i}^{-1},$
@@
$X_{T,j}=\prod_{i=1}^{j}\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)}\times_{B}\prod_{i=j+1}^{r}\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\bigr{)},$
@@
The $\mathbb{T}$-fixed loci of $X_{T,\pm}$ are indexed by (not necessarily injective) functions $f\colon\{1,\dots,r\}\to\{1,\dots,n\}$. For such an $f$, the associated fixed locus $B_{f}$ is the subspace of $X_{T,+}$ such that, on the $i$th factor of $\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)}$ in the fiber product, the homogeneous coordinates are all zero except for in the $f(i)$th factor of $\mathbb{P}(E)$. In other words,
$B\cong B_{f}=\mathbb{P}\bigl{(}L_{f(1)}\bigr{)}\times_{B}\cdots\times_{B}\mathbb{P}\bigl{(}L_{f(r)}\bigr{)}\subset\mathbb{P}(E)\times_{B}\cdots\times_{B}\mathbb{P}(E)\subset X_{T,+}.$
@@
We will make use of the following weak form of the abelian/non-abelian correspondence in cohomology.
@@
$H^{*}_{\mathbb{T}}(X_{\pm})\to\bigl{(}H^{*}_{\mathbb{T}}\bigl{(}V^{s,\pm}(G)/T\bigr{)}\bigr{)}^{W}$
is an isomorphism, where $V^{s,\pm}(G)$ denotes the stable locus of $V$ with respect to $\det^{\pm}\colon G\to\mathbb{C}^{*}$ and $W$ denotes the Weyl group. Furthermore, the pullback map
$H^{*}_{\mathbb{T}}(X_{T,\pm})\to H^{*}_{\mathbb{T}}\bigl{(}V^{s,\pm}(G)/T\bigr{)}$
@@
See [6, Proposition 1] for details. The proposition allows us to lift cohomology classes in $X_{\pm}$ to $W$-invariant classes in $H^{*}_{\mathbb{T}}(X_{T,\pm})$.
For future use we will use $F_{\pm}$ to denote the set of functions $f\colon\{1,\dots,r\}\to\{1,\dots,n\}$, i.e., $F_{\pm}$ indexes the set of fixed loci of $X_{T,\pm}$. We will use ${\rm In}_{\pm}$ to denote the set of injective functions $\{1,\dots,r\}\to\{1,\dots,n\}$, noting that ${\rm In}_{\pm}$ indexes the fixed loci of $V^{s,\pm}(G)/T.$
2.3 Fourier–Mukai transform
We can define a resolution of the birational map $f\colon X_{-}\dashrightarrow X_{+}$ as follows. Let
$\tilde{V}=\mathcal{H}{\rm om}(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$
and let $\tilde{G}={\rm GL}_{r}\times{\rm GL}_{r}$. Consider the right action of $\tilde{G}$ on $\tilde{V}$ given by
$\bigl{(}\tilde{M},\tilde{D},\tilde{N}\bigr{)}\cdot(g_{1},g_{2})=\bigl{(}g_{1}^{-1}\tilde{M},g_{2}^{-1}\tilde{D}g_{1},\tilde{N}g_{2}\bigr{)}.$
Let $\tilde{\theta}\in\operatorname{Hom}\bigl{(}\tilde{G},\mathbb{C}^{*}\bigr{)}$ denote the character